OpenStudy (anonymous):

In the equation x – 4y = 2, for what value of x will y = 0? A. 2 B. 6 C. –2 D. 0

OpenStudy (uri):

x – 4y = 2 x-4(0)=2 x-0=2 x=2+0 x=2